# Exploring the Frontier of Vision-Language Models: A Survey of Current Methodologies and Future Directions

### v1

@@ 2.1 Vision-Language Understanding
LLaVA
90.92
50.37 R/A
41.3
81.9
23.87B
109113B
LaVIN 13B
90.83
LaVIN 7B
89.41
BLIP-2
61img
19.6
71.6
85.3V13B
41
41
38.1w
22.4
InstructBLIPV13B
63.1
34.5
82.8
88.57 R/A
49.5
58.2w
25.6
InstructBLIPV7B
60.5
49.2
60.9w
26.2
964
MetaCLIP
68.77
MiniGPT4
58.7
79.67 R/A
30.8
24.4
900
Qwen-VL
67.1img
35.2
81
78.8
59.3
Qwen-VL-Chat
68.2img
38.9
85.8
78.2
57.5
GPT-4V
85
23
15.2(S)
0VQAS
10EM
78.5
1349
LLaVa 1.5 V7B
66.8
50
85.9
78.5
62
63.4w
LLaVa 1.5V13B
71.6
53.6
85.9
80
63.3
70.7
BakLLaVa-1
66.7
86.6
LLaMa-VIDV13B
70
54.3
86
80
65
LLaMa-VID V7B
68.3
54.2
86
79.3
64.3
Sphinx
Gemini
80.8u
IDEFICS-9B
35.5
50.9
38.4
997
IDEFICS-80B
36
60
45.2
Flamingo-9B
28.8
61.5
51.8
TinyGPT
33.4
33.6
Flamingo-80B
31.6
56.3
KOSMOS-1
29.2
65.2
51
Ferret-7B
80.39
90.24 R/A
86.7
Ferret-13B
81.13
87.5
mPLUG-Owl
53.97 R/A
1025
KOSMOS-2
80.5
51.1
70
Fuyu-8B
74.1
42.1
LLaMa-Adapter-v2-7b
31.4
1066
PaLM-E12B
76.2
PaLM-E562B
80
Frozen
29.5
Emu237B
33.3
Emu2-Chat
54.9
84.9
65.1
LyricsV13B
71.1img
37.6
8534
81.2
62.4
Table 1: Table showing the comparative analysis of multiple Visual Language Models on 10 benchmark datasets: Science-QA Lu et al. (2022), VizWiz Danna et al. (2018), Flickr30K A et al. (2016), POPE Fan et al. (2023), VQAv2 Goyal et al. (2017), GQA Hudson and Manning (2019), LLaVaBench Liu et al. (2023c), Chart-QA Masry et al. (2022), $M$-Vet Yu et al. (2023), ViSiTBench Bitton et al. (2023) . V: Vicuna, , img: only image, R/A: Random and Accuracy has been reported, VQAS: VQA Score, EM: EM Accuracy, w: in-the-wild version, u: Ultra version.XB: The model is of X billion parameters.
@@ 2.2 Text generation with Multimodal Input
GPT-4V OpenAI (2023): GPT-4V marks a significant advancement, allowing users to instruct GPT-4 to analyze image inputs. OpenAI conducted extensive safety evaluations and preparations for GPT-4V, building on the safety work done for GPT-4. The training process involved predicting the next word in a document, utilizing a vast dataset of text and image data. GPT-4V inherits both text and vision capabilities, presenting novel features at their intersection. The system card outlines OpenAI’s preparation, early access period, safety evaluations, red team assessments, and implemented mitigations before broad release.
@@
LLaVA Liu et al. (2023a): The passage describes LLaVA, an open-source multimodal framework designed to enhance LLMs for understanding both language and images. It utilizes language-only GPT-4 to generate data for instruction-following tasks in a multimodal context. LLaVA integrates a vision encoder from CLIP with the LLM, enabling it to process visual information alongside language. The model undergoes pre-training on image-text pairs and fine-tuning for end-to-end multimodal understanding, resulting in a versatile multimodal chatbot. Early experiments show LLaVA’s impressive performance, achieving an 85.1% relative score compared to GPT-4 on a synthetic instruction-following dataset and setting a new state-of-the-art accuracy of 92.53% on Science QA tasks when combined with GPT-4.
@@
Flamingo Alayrac et al. (2022): Flamingo introduces novel architectural features to seamlessly integrate vision-only and language-only models. By incorporating interleaved cross-attention layers with frozen language-only self-attention layers, Flamingo excels in handling sequences of intermixed visual and textual data. It adopts a Perceiver-based architecture to convert input sequence data, like videos, into a fixed set of visual tokens. Leveraging large-scale multimodal web corpora with interleaved text and images, Flamingo demonstrates remarkable few-shot learning capabilities across various benchmarks, surpassing models fine-tuned on significantly more task-specific data. This showcases its adaptability and efficiency in rapidly adapting to diverse image and video understanding tasks with limited examples.OpenFlamingo is an ongoing initiative aiming to create an open-source version of DeepMind’s Flamingo models. Across seven vision-language datasets, OpenFlamingo models consistently achieve performance levels ranging from 80% to 89% of their counterparts in the Flamingo models. Med-Flamingo Moor et al. (2023), a medical-focused multimodal few-shot learner based on OpenFlamingo-9B, achieves up to a 20% improvement in generative medical Visual Question Answering. It pioneers human evaluation in this context, involving clinicians in an interactive assessment, and enables applications like rationale generation.
@@
PALM-E Driess et al. (2023): PALM-E emerges as an innovative Embodied Multimodal Language Model, meticulously crafted to navigate real-world scenarios by fusing language comprehension with continuous sensor inputs. This model is the result of a collaborative endeavor between TU Berlin and Google Research, signifying a pivotal advancement in the realm of multimodal AI. It incorporates real-world continuous sensor modalities into language models, training multi-modal sentences end-to-end with a pre-trained large language model. The model successfully tackles multiple embodied tasks, including robotic manipulation planning, visual question answering, and captioning. PaLM-E, the largest model with 562B parameters, demonstrates state-of-the-art performance in both embodied reasoning tasks and visual-language domains like OK-VQA. The model operates on multimodal sentences, showcasing positive transfer of knowledge from visual-language domains to embodied reasoning tasks, emphasizing its versatility and scalability. PaLM-E faces limitations in relying on low-level language-conditioned policies for robotic tasks, prompting the proposal of self-supervised entity-centric labeling to enhance guidance in intricate tasks.
@@
BLIP Li et al. (2023a): BLIP is a novel Vision-Language Pre-training (VLP) framework, that surpasses existing models, excelling in both understanding and generation tasks, addressing noisy training limitations.BLIP introduces the Multimodal Mixture of Encoder-Decoder (MED) architecture, utilizing Image-Text Contrastive (ITC), Image-Text Matching (ITM), and Language Modeling (LM) objectives for pre-training. Captioning and Filtering (CapFilt) enhances data quality, improving downstream task performance.BLIP, implemented in PyTorch and pre-trained on a diverse 14 million image dataset, demonstrated notable improvements in downstream tasks like image-text retrieval and captioning. Leveraging nucleus sampling and effective parameter sharing, BLIP outperformed existing models on standard datasets.
@@
BLIP-2 Li et al. (2023a): It introduces a cost-effective vision-language pre-training strategy with frozen image encoders and large language models. The Querying Transformer achieves state-of-the-art performance in vision-language tasks with fewer parameters, addressing interoperability challenges between different modality embeddings. BLIP-2 introduced a fresh component known as the Querying Transformer (Q-Former), serving as a trainable bridge between a static image encoder and a static LLM. Illustrated in Figure[3], the Q-Former undergoes a two-stage pre-training regimen. Initially, the focus lies on vision-language representation learning, compelling the Q-Former to grasp visual elements most pertinent to the accompanying text. Subsequently, the emphasis shifts to vision-to-language generative learning, linking the Q-Former’s output to a fixed LLM and refining its capabilities so that the visual representations it produces can be effectively interpreted by the LLM.
@@
InstructBLIP Dai et al. (2023): InstructBLIP employs instruction-aware visual feature extraction, enhancing its ability to extract informative features tailored to provided instructions. Achieving state-of-the-art zero-shot performance across 13 held-out datasets, InstructBLIP outperforms BLIP-2 and larger models like Flamingo. The models also excel in downstream tasks, demonstrating a 90.7% accuracy on ScienceQA IMG Lu et al. (2022), and showcase qualitative advantages over concurrent multimodal models in diverse capabilities such as visual scene understanding, knowledge-grounded image description, and multi-turn visual conversation.
@@
KOSMOS-1 Huang et al. (2023): KOSMOS-1 is a VLM from Microsoft. Trained on a web-scale multimodal corpus, KOSMOS-1 excels in language understanding and generation, OCR-free NLP, and various perception-language tasks, showcasing its capabilities in image captioning and visual question answering. Using a Transformer-based architecture, KOSMOS-1 aligns vision with large language models. Its training involves diverse multimodal corpora, including The Pile and Common Crawl, with language-only instruction tuning. Additionally, the model excels in chain-of-thought prompting, generating rationales before addressing complex question-answering tasks. Overall, KOSMOS-1 represents a significant advancement in the field of Multimodal Large Language Models, offering robust performance across a wide range of tasks.
@@
KOSMOS-2 Peng et al. (2023): KOSMOS-2 again from Microsoft Research, advances traditional models by introducing capabilities for perceiving object descriptions, such as bounding boxes, and grounding text in the visual world. Utilizing a unique representation format for referring expressions, KOSMOS-2 links text spans to spatial locations in images. Employing a sophisticated image processing approach, the model combines vision encodings with location tokens to understand and relate specific image areas to textual descriptions. Built on the KOSMOS-1 architecture, this Transformer-based causal language model represents a substantial step towards Embodiment AI and the convergence of language, multimodal perception, action, and world modeling, showcasing its potential for advancing artificial general intelligence.
@@
MultiInstruct Xu et al. (2022): MultiInstruct presents a benchmark dataset for multimodal instruction tuning, featuring 62 tasks across 10 categories. Utilizing the OFA pre-trained multimodal language model, the study focuses on enhancing zero-shot performance on diverse tasks through large-scale text-only instruction datasets like Natural Instructions. Results show strong zero-shot performance and reduced model sensitivity to instruction variations. Comparative analysis of transfer learning strategies indicates improved robustness across multimodal tasks. Increasing task clusters during training enhances overall performance, supporting the effectiveness of MultiInstruct.
@@
IDEFICS Laurençon et al. (2023) : IDEFICS, an open-access reproduction of the closed-source vision-language model Flamingo by DeepMind, boasts 80 billion parameters and is available on HuggingFace. It performs well in image-text benchmarks, such as visual question answering and image captioning, utilizing in-context few-shot learning. IDEFICS has two versions – an 80 billion parameters model and a 9 billion parameters model.
@@
PaLI Chen et al. (2022): PALI, or Pathways Language and Image model, from Google Research, leverages large pre-trained encoder-decoder language models and vision transformers for joint language and vision modeling. The model achieves state-of-the-art results in various vision and language tasks across 100+ languages by utilizing a diverse multilingual dataset containing 10B images and texts. With a simple, modular, and scalable design, PaLI highlights the importance of joint scaling in vision and language components for effective training and performance.
@@
Frozen Tsimpoukelli et al. (2021): Frozen, a multimodal few-shot learning approach developed by DeepMind, leveraging a pre-trained, frozen language model and a vision encoder trained on aligned image and caption data to enable rapid learning of new tasks with just a few examples, demonstrating its effectiveness in tasks such as learning words for new objects, visual question-answering, and incorporating outside knowledge across various benchmarks. The approach utilizes gradients through a frozen language model’s self-attention layers to train the vision encoder. The system’s notable limitation lies in its suboptimal performance on tasks learned with few shots compared to state-of-the-art models using full training sets, highlighting the potential for enhanced zero-shot and few-shot generalization through further improvements in accuracy and reduced seed requirements.
@@
Qwen-VL Bai et al. (2023a) : The Qwen-VL series, introduced as large-scale vision-language models, encompasses Qwen-VL and Qwen-VL-Chat, demonstrating exceptional performance in tasks such as image captioning, question answering, visual localization, and versatile interactions. Qwen-VLs demonstrate outstanding performance across various vision-centric tasks, surpassing counterparts of similar scale. Their exceptional accuracy extends beyond traditional benchmarks like captioning and question-answering to include recent dialogue benchmarks. Trained on multilingual image-text data, with a significant portion in English and Chinese, Qwen-VLs naturally support multiple languages. They handle multiple images concurrently during training, enabling Qwen-Chat-VL to contextualize and analyze complex scenarios. With higher-resolution inputs and fine-grained training data, Qwen-VLs excel in fine-grained visual understanding, outperforming existing vision-language models in grounding, text comprehension, question answering, and dialogue tasks.
@@
Fuyu-8B Bavishi et al. (2023): Fuyu-8B, a multi-modal text and image transformer developed by Adept AI, offers a simplified yet powerful solution tailored for digital agents. Its straightforward architecture and training process enhance comprehension, scalability, and deployment, making it ideal for various applications. Specifically designed for digital agents, Fuyu-8B seamlessly handles arbitrary image resolutions and excels in tasks like graph and diagram comprehension, UI-based queries, and rapid processing of large images within 100 milliseconds. Despite its optimization for Adept’s use case, Fuyu-8B delivers impressive performance in standard image understanding benchmarks such as visual question-answering and natural image captioning. Architecturally, Fuyu adopts a vanilla decoder-only transformer, efficiently processing image patches by linear projection into the first layer. Its versatility in supporting diverse image resolutions is achieved by treating image tokens like text tokens, utilizing raster-scan order, and signaling line breaks for adaptability.
@@
Sphinx Lin et al. (2023a): SPHINX is a versatile VLM that integrates model weights, tuning tasks, and visual embeddings to enhance its capabilities. It unfreezes the large language model during pre-training to strengthen vision-language alignment and efficiently mixes weights from LLMs trained on real-world and synthetic data for robust understanding. By incorporating diverse tasks like region-level understanding and human pose estimation, SPHINX achieves mutual enhancement across different scenarios. It also extracts comprehensive visual embeddings from various sources, enriching language models with robust image representations. SPHINX demonstrates superior multi-modal understanding across applications and introduces an efficient strategy for capturing fine-grained details in high-resolution images, excelling in visual parsing and reasoning tasks.
@@
Mirasol Piergiovanni et al. (2023): Mirasol, from Google DeepMind and Google Research, is a multimodal autoregressive model designed to handle both time-aligned modalities (audio, video) and non-aligned modality (text). The architecture involves segmenting long video-audio sequences into manageable chunks, passing them through respective encoders, and using a Combiner to fuse video and audio features. Autoregressive training predicts sequential features, with a separate Transformer block integrating textual prompts through cross-modal attention. This enables enriched contextual understanding, showcasing a comprehensive approach to multimodal learning and generation. Pretrained on 12% of VTP, the model uniformly weighed losses in pretraining, with a tenfold emphasis on unaligned text loss during fine-tuning. Ablation studies underscore its ability to maintain content consistency and adapt to dynamic changes in video-audio sequence
@@
MiniGPT-4 Zhu et al. (2023a): MiniGPT-4 combines a frozen visual encoder (ViT + Q-Former from BLIP-2) with LLM using a single trainable projection layer. Pretrained on aligned image-text pairs and fine-tuned on detailed image descriptions, MiniGPT-4 exhibits GPT-4-like capabilities without training vision or language modules separately. The finetuning process enhances language outputs, demonstrating diverse skills like meme interpretation, recipe generation, and poem composition. The model’s architecture involves a vision encoder, linear projection layer, and large language model.
@@
MiniGPT-v2 Chen et al. (2023d): The model architecture of MiniGPT-v2 consists of a ViT visual backbone, a linear projection layer, and a large language model like LLaMA-2 Touvron et al. (2023). The ViT backbone is frozen during training, and four adjacent visual output tokens are concatenated and projected into LLaMA-2 space. Task-specific identifiers are incorporated during training using a three-stage strategy with weakly labeled image-text datasets and multi-modal instructional datasets. The model demonstrates superior performance in visual question-answering and visual grounding, outperforming other generalist models. The use of task identifier tokens enhances efficiency in multi-task learning, contributing to its state-of-the-art performance. Challenges include occasional hallucinations, emphasizing the need for more high-quality image-text-aligned data.
@@
LLaVA-Plus Liu et al. (2023e): LLaVA-Plus, is a general-purpose multimodal assistant designed to enhance LMMs through visual instruction tuning. The model maintains a skill repository with diverse vision and vision-language pre-trained models, activating relevant tools in response to user inputs for various tasks. Trained on multimodal instruction-following data, LLaVA-Plus covers tool usage in visual understanding, generation, and external knowledge retrieval, surpassing its predecessor, LLaVA, in both existing and new capabilities.The training approach involves using GPT-4 to generate instruction data and integrating new tools through instruction tuning, enabling continuous enhancement. LLaVA-Plus demonstrates state-of-the-art performance on VisiT-Bench, a real-life multimodal task benchmark, excelling in tool use compared to other tool-augmented LLMs.
BakLLaVA Bitton et al. (2023): It is a VLM developed by LAION, Ontocord, and Skunkworks AI. BakLLaVA uses a Mistral 7B base augmented with the LLaVA 1.5 architecture. Used in combination with llama.cpp. BakLLaVA is a faster and less resource-intensive alternative to GPT-4 with Vision.

### v2

@@ 2.1 Vision-Language Understanding
LLaVA 90.92 - - 50.37 R/A - 41.3 81.9 - 23.87B 109113B
LaVIN 13B 90.83 - - - - - - - - -
LaVIN 7B 89.41 - - - - - - - - -
BLIP-2 61img 19.6 71.6 85.3V13B 41 41 38.1w - 22.4 -
InstructBLIPV13B 63.1 34.5 82.8 88.57 R/A - 49.5 58.2w - 25.6 -
InstructBLIPV7B 60.5 - - - - 49.2 60.9w - 26.2 964
MetaCLIP 68.77 - - - - - - - - -
MiniGPT4 58.7 - - 79.67 R/A - 30.8 - - 24.4 900
Qwen-VL 67.1img 35.2 81 - 78.8 59.3 - - - -
Qwen-VL-Chat 68.2img 38.9 85.8 - 78.2 57.5 - - - -
GPT-4V 85 23 15.2(S) - 0VQAS 10EM - 78.5 - 1349
LLaVa 1.5 V7B 66.8 50 - 85.9 78.5 62 63.4w - - -
LLaVa 1.5V13B 71.6 53.6 - 85.9 80 63.3 70.7 - - -
BakLLaVa-1 66.7 - - 86.6 - - - - - -
LLaMa-VIDV13B 70 54.3 - 86 80 65 - - - -
LLaMa-VID V7B 68.3 54.2 - 86 79.3 64.3 - - - -
Sphinx - - - - - - - - - -
Gemini - - - - - - - 80.8u - -
IDEFICS-9B - 35.5 - - 50.9 38.4 - - - 997
IDEFICS-80B - 36 - - 60 45.2 - - - -
Flamingo-9B - 28.8 61.5 - 51.8 - - - - -
TinyGPT - 33.4 - - - 33.6 - - - -
Flamingo-80B - 31.6 - - 56.3 - - - - -
KOSMOS-1 - 29.2 65.2 - 51 - - - - -
Ferret-7B - - 80.39 90.24 R/A - - 86.7 - - -
Ferret-13B - - 81.13 - - - 87.5 - - -
mPLUG-Owl - - - 53.97 R/A - - - - - 1025
KOSMOS-2 - - 80.5 - 51.1 - 70 - - -
Fuyu-8B - - - 74.1 - - - 42.1 - -
LLaMa-Adapter-v2-7b - - - - - - - - 31.4 1066
PaLM-E12B - - - - 76.2 - - - - -
PaLM-E562B - - - - 80 - - - - -
Frozen - - - - 29.5 - - - - -
Emu237B - - - - 33.3 - - - - -
Emu2-Chat - 54.9 - - 84.9 65.1 - - - -
LyricsV13B 71.1img 37.6 8534 - 81.2 62.4 - - - -
Table 1: Table showing the comparative analysis of multiple Visual Language Models on 10 benchmark datasets: Science-QA [57], VizWiz [24], Flickr30K [2], POPE [27], VQAv2 [32], GQA [36], LLaVaBench [53], Chart-QA [62], $M$-Vet [105], ViSiTBench [10] . V: Vicuna, , img: only image, R/A: Random and Accuracy have been reported, VQAS: VQA Score, EM: EM Accuracy, w: in-the-wild version, u: Ultra version.XB: The model is of X billion parameters.
@@ 2.2 Text generation with Multimodal Input
GPT-4V [64]: GPT-4V marks a significant advancement, allowing users to instruct GPT-4 to analyze image inputs. OpenAI conducted extensive safety evaluations and preparations for GPT-4V, building on the safety work done for GPT-4. The training process involved predicting the next word in a document, utilizing a vast dataset of text and image data. GPT-4V inherits both text and vision capabilities, presenting novel features at their intersection. The system card outlines OpenAI’s preparation, early access period, safety evaluations, red team assessments, and implemented mitigations before broad release.
@@
LLaVA [51]: LLaVA, an open-source multimodal framework designed to enhance LLMs for understanding both language and images. It utilizes language-only GPT-4 to generate data for instruction-following tasks in a multimodal context. LLaVA integrates a vision encoder from CLIP with the LLM, enabling it to process visual information alongside language. The model undergoes pre-training on image-text pairs and fine-tuning for end-to-end multimodal understanding, resulting in a versatile multimodal chatbot. LLaVA demonstrates impressive multimodal chat abilities and achieves an 85.1% relative score compared to GPT-4 on a synthetic multimodal instruction-following dataset. Upon fine-tuning on the Science QA dataset, LLaVA and GPT-4 together achieve a new state-of-the-art accuracy of 92.53%.
@@
Flamingo [4]: Flamingo introduces novel architectural features to seamlessly integrate vision-only and language-only models. By incorporating interleaved cross-attention layers with frozen language-only self-attention layers, Flamingo excels in handling sequences of intermixed visual and textual data. It adopts a Perceiver-based architecture to convert input sequence data, like videos, into a fixed set of visual tokens. Leveraging large-scale multimodal web corpora with interleaved text and images, Flamingo demonstrates remarkable few-shot learning capabilities across various benchmarks, surpassing models fine-tuned on significantly more task-specific data. This showcases its adaptability and efficiency in rapidly adapting to diverse image and video understanding tasks with limited examples.The OpenFlamingo project is an ongoing effort dedicated to crafting an open-source rendition of DeepMind’s Flamingo models. Across a spectrum of seven vision-language datasets, the performance of OpenFlamingo models consistently ranges between 80% and 89% when compared to the original Flamingo models. Med-Flamingo [63], a medical-focused multimodal few-shot learner based on OpenFlamingo-9B, achieves up to a 20% improvement in generative medical Visual Question Answering. It pioneers human evaluation in this context, involving clinicians in an interactive assessment, and enables applications like rationale generation.
@@
PALM-E [25]: PALM-E emerges as an innovative Embodied Multimodal Language Model, meticulously crafted to navigate real-world scenarios by fusing language comprehension with continuous sensor inputs. This model is the result of a collaborative endeavor between TU Berlin and Google Research, signifying a pivotal advancement in the realm of multimodal AI. By integrating continuous sensor modalities from the real world into language models, this approach enables end-to-end training of multimodal sentences using a pre-trained large language model. It effectively addresses various embodied tasks such as robotic manipulation planning, visual question answering, and captioning. PaLM-E, the most extensive model boasting 562 billion parameters, exhibits cutting-edge performance across embodied reasoning tasks and visual-language domains like OK-VQA. Operating on multimodal sentences, it demonstrates the transfer of knowledge from visual-language domains to embodied reasoning tasks, underscoring its adaptability and scalability. PaLM-E faces limitations in relying on low-level language-conditioned policies for robotic tasks, prompting the proposal of self-supervised entity-centric labeling to enhance guidance in intricate tasks.
@@
BLIP [42]: BLIP stands out as an innovative Vision-Language Pre-training framework, overcoming challenges associated with noisy training data. Central to BLIP is its Multimodal Mixture of Encoder-Decoder (MED) architecture, which incorporates Image-Text Contrastive (ITC),Language Modeling (LM) and Image-Text Matching (ITM) objectives during pre-training. Captioning and Filtering (CapFilt) enhances data quality, improving downstream task performance.BLIP, implemented in PyTorch and pre-trained on a diverse 14 million image dataset, demonstrated notable improvements in downstream tasks like image-text retrieval and captioning. Leveraging nucleus sampling and effective parameter sharing, BLIP outperformed existing models on standard datasets.
@@
BLIP-2 [42]:BLIP-2 introduces an efficient strategy for vision-language pre-training, utilizing frozen image encoders alongside large language models. The Querying Transformer within BLIP-2 achieves top-tier performance in vision-language tasks while employing fewer parameters, effectively addressing challenges in interoperability among different modality embeddings. A novel addition brought by BLIP-2 is the Querying Transformer (Q-Former), which acts as a trainable link between static image encoders and a fixed LLM. Depicted in Figure[3], the Q-Former undergoes a two-stage pre-training process. Initially, it focuses on learning representations that bridge vision and language, enabling it to comprehend visual elements crucial to accompanying text. Later, the emphasis shifts to generative learning, connecting the Q-Former’s output to the fixed LLM and refining its ability to generate visual representations interpretable by the LLM.
@@
InstructBLIP [23]: InstructBLIP employs instruction-aware visual feature extraction, enhancing its ability to extract informative features tailored to provided instructions. Achieving state-of-the-art zero-shot performance across 13 held-out datasets, InstructBLIP outperforms BLIP-2 and larger models like Flamingo. The models also excel in downstream tasks, demonstrating a 90.7% accuracy on ScienceQA IMG [57], and showcase qualitative advantages over concurrent multimodal models in diverse capabilities such as knowledge-grounded image description, visual scene understanding, and multi-turn visual conversation.
@@
KOSMOS-1 [35]: KOSMOS-1 is a VLM from Microsoft. Trained on a web-scale multimodal corpus, KOSMOS-1 excels in language understanding and generation, OCR-free NLP, and various perception-language tasks, showcasing its capabilities in image captioning and visual question answering. Using a Transformer-based architecture, KOSMOS-1 aligns vision with large language models. Its training involves diverse multimodal corpora, including The Pile and Common Crawl, with language-only instruction tuning. Additionally, the model excels in chain-of-thought prompting, generating rationales before addressing complex question-answering tasks. Overall, KOSMOS-1 represents a significant advancement in the field of Multimodal Large Language Models, offering robust performance across a wide range of tasks.
@@
KOSMOS-2 [67]: KOSMOS-2 again from Microsoft Research, advances traditional models by introducing capabilities for perceiving object descriptions, such as bounding boxes, and grounding text in the visual world. Utilizing a unique representation format for referring expressions, KOSMOS-2 links text spans to spatial locations in images. Employing a sophisticated image processing approach, the model combines vision encodings with location tokens to understand and relate specific image areas to textual descriptions. Constructed upon the foundational KOSMOS-1 architecture, this causal language model, based on Transformers, signifies a significant advancement towards Embodiment AI. It marks a pivotal stride towards the integration of language, visual perception, action, and world modeling, bringing them closer together in convergence for advancing artificial general intelligence.
@@
MultiInstruct [95]: MultiInstruct presents a benchmark dataset for multimodal instruction tuning, featuring 62 tasks across 10 categories. Utilizing the OFA pre-trained multimodal language model, the study focuses on enhancing zero-shot performance on diverse tasks through large-scale text-only instruction datasets like Natural Instructions. Results show strong zero-shot performance and reduced model sensitivity to instruction variations. Comparative analysis of transfer learning strategies indicates improved robustness across multimodal tasks. Increasing task clusters during training enhances overall performance, supporting the effectiveness of MultiInstruct.
@@
IDEFICS [39] : IDEFICS, an open-access reproduction of the closed-source vision-language model Flamingo by DeepMind, boasts 80 billion parameters and is available on HuggingFace. It performs well in image-text benchmarks, such as visual question answering and image captioning, utilizing in-context few-shot learning. IDEFICS has two versions – an 80 billion parameters model and a 9 billion parameters model.
@@
PaLI [12]: PALI, or Pathways Language and Image model, from Google Research, leverages large pre-trained encoder-decoder language models and vision transformers for joint language and vision modeling. The model achieves state-of-the-art results in various vision and language tasks across 100+ languages by utilizing a diverse multilingual dataset containing 10B images and texts. With a simple, modular, and scalable design, PaLI highlights the importance of joint scaling in vision and language components for effective training and performance.
@@
Frozen [83]: Frozen, an innovative multimodal learning approach crafted by DeepMind, combines vision encoders trained on image captioning data with a frozen language model. This design empowers the model to swiftly adapt to new tasks in a few-shot setup, showcasing its efficacy across a spectrum of challenges such as visual question answering across diverse benchmark datasets. This approach trains the vision encoder by backpropagating the gradients thorough the frozen language model’s self-attention layers.The system’s notable limitation lies in its suboptimal performance on tasks learned with few shots compared to state-of-the-art models using full training sets, highlighting the potential for enhanced zero-shot and few-shot generalization through further improvements in accuracy and reduced seed requirements.
@@
Qwen-VL [5] : Qwen-VL series, introduced as large vision-language models, encompasses Qwen-VL and Qwen-VL-Chat, demonstrating exceptional performance in tasks such as image captioning, question answering, visual localization, and versatile interactions. Qwen-VLs demonstrate outstanding performance across various vision-centric tasks, surpassing counterparts of similar scale. Their exceptional accuracy extends beyond traditional benchmarks like captioning and question-answering to include recent dialogue benchmarks. Trained on multilingual image-text data, with a significant portion in English and Chinese, Qwen-VLs naturally support multiple languages. They handle multiple images concurrently during training, enabling Qwen-Chat-VL to contextualize and analyze complex scenarios. With higher-resolution inputs and fine-grained training data, Qwen-VLs excel in fine-grained visual understanding, outperforming existing vision-language models in grounding, text comprehension, question answering, and dialogue tasks.
@@
Fuyu-8B [9]: Fuyu-8B, a multi-modal text and image transformer developed by Adept AI, offers a simplified yet powerful solution tailored for digital agents. Its straightforward architecture and training process enhance comprehension, scalability, and deployment, making it ideal for various applications. Specifically designed for digital agents, Fuyu-8B seamlessly handles arbitrary image resolutions and excels in tasks like graph and diagram comprehension, UI-based queries, and rapid processing of large images within 100 milliseconds. Despite its optimization for Adept’s use case, Fuyu-8B delivers impressive performance in standard image understanding benchmarks such as visual question-answering and natural image captioning. Architecturally, Fuyu adopts a vanilla decoder-only transformer, efficiently processing image patches by linear projection into the first layer. Its versatility in supporting diverse image resolutions is achieved by treating image tokens like text tokens, utilizing raster-scan order, and signaling line breaks for adaptability.
@@
Sphinx [48]: SPHINX is a versatile VLM that integrates model weights, tuning tasks, and visual embeddings to enhance its capabilities. It unfreezes the large language model during pre-training to strengthen vision-language alignment and efficiently mixes weights from LLMs trained on real-world and synthetic data for robust understanding. By incorporating diverse tasks like region-level understanding and human pose estimation, SPHINX achieves mutual enhancement across different scenarios. It also extracts comprehensive visual embeddings from various sources, enriching language models with robust image representations. SPHINX demonstrates superior multi-modal understanding across applications and introduces an efficient strategy for capturing fine-grained details in high-resolution images, excelling in visual parsing and reasoning tasks.
@@
Mirasol [68]: Mirasol, from Google DeepMind and Google Research, is a multimodal autoregressive model designed to handle both time-aligned modalities (audio, video) and non-aligned modality (text). The architecture involves segmenting long video-audio sequences into manageable chunks, passing them through respective encoders, and using a Combiner to fuse video and audio features. Autoregressive training predicts sequential features, with a separate Transformer block integrating textual prompts through cross-modal attention. This enables enriched contextual understanding, showcasing a comprehensive approach to multimodal learning and generation. Pretrained on 12% of VTP, the model uniformly weighed losses in pretraining, with a tenfold emphasis on unaligned text loss during fine-tuning. Ablation studies underscore its ability to maintain content consistency and adapt to dynamic changes in video-audio sequence
@@
MiniGPT-4 [119]: MiniGPT-4 combines a frozen visual encoder (ViT + Q-Former from BLIP-2) with LLM using a single trainable projection layer. Pretrained on aligned image-text pairs and fine-tuned on detailed image descriptions, MiniGPT-4 exhibits GPT-4-like capabilities without training vision or language modules separately. The finetuning process enhances language outputs, demonstrating diverse skills like meme interpretation, recipe generation, and poem composition. The model’s architecture involves a vision encoder, linear projection layer, and large language model.
@@
MiniGPT-v2 [15]: The model architecture of MiniGPT-v2 consists of a ViT visual backbone, a projection layer for dimension matching , and a large language model like LLaMA-2 [82] for the final generation . The ViT backbone is frozen during training, and four adjacent visual output tokens are concatenated and projected into LLaMA-2 space. Task-specific identifiers are incorporated during training using a three-stage strategy with weakly labeled image-text datasets and multi-modal instructional datasets. The model demonstrates superior performance in visual question-answering and visual grounding, outperforming other generalist models. The use of task identifier tokens enhances efficiency in multi-task learning, contributing to its state-of-the-art performance. Challenges include occasional hallucinations, emphasizing the need for more high-quality image-text-aligned data.
@@
LLaVA-Plus [55]: LLaVA-Plus, is a general-purpose multimodal assistant designed to enhance LMMs through visual instruction tuning. The model maintains a skill repository with diverse vision and vision-language pre-trained models, activating relevant tools in response to user inputs for various tasks. Trained on multimodal instruction-following data, LLaVA-Plus covers tool usage in visual understanding, generation, and external knowledge retrieval, surpassing its predecessor, LLaVA, in both existing and new capabilities.The training approach involves using GPT-4 to generate instruction data and integrating new tools through instruction tuning, enabling continuous enhancement. LLaVA-Plus demonstrates state-of-the-art performance on VisiT-Bench, a real-life multimodal task benchmark, excelling in tool use compared to other tool-augmented LLMs.
BakLLaVA [10]: BakLLaVA represents a Visual Language Model (VLM) crafted through a collaborative endeavor involving LAION, Ontocord, and Skunkworks AI. It harnesses the power of the Mistral 7B base, enhanced by the innovative LLaVA 1.5 architecture. When paired with the llama.cpp framework, BakLLaVA emerges as a swifter and more resource-efficient alternative to GPT-4 with Vision capabilities.